OpenStudy (anonymous):

3/8(5+6x)+3>21 the solution is

OpenStudy (anonymous):

subtract 3 multiply by 8/3 subtract 5 divide by 6

OpenStudy (anonymous):

would you like me to write it out?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sir see my question of matrices i posted

OpenStudy (anonymous):

\[\frac{3}{8}(5+6x)+3>21\] \[\frac{3}{8}(5+6x)>18\] \[5+6x>18\times \frac{8}{3}\] \[5+6x>48\] \[6x>43\] \[x>\frac{43}{6}\]